Single Plantwide and Multiple Production Department Factory Overhead Rate Methods and Product Cost Distortion Eclipse Motor...

  1. Single Plantwide and Multiple Production Department Factory Overhead Rate Methods and Product Cost Distortion

    Eclipse Motor Company manufactures two types of specialty electric motors, a commercial motor and a residential motor, through two production departments, Assembly and Testing. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering using the multiple production department factory overhead rate method. The following factory overhead was budgeted for Eclipse:

    Assembly Department $280,000
    Testing Department 800,000
    Total $1,080,000

    Direct machine hours were estimated as follows:

    Assembly Department 4,000 hours
    Testing Department 5,000
    Total 9,000 hours

    In addition, the direct machine hours (dmh) used to produce a unit of each product in each department were determined from engineering records, as follows:

    Commercial Residential
    Assembly Department 2.0 dmh 3.0 dmh
    Testing Department 6.0 1.5
    Total machine hours per unit 8.0 dmh 4.5 dmh

    a. Determine the per-unit factory overhead allocated to the commercial and residential motors under the single plantwide factory overhead rate method, using direct machine hours as the allocation base.

    Commercial $ per unit
    Residential $ per unit

    b. Determine the per-unit factory overhead allocated to the commercial and residential motors under the multiple production department factory overhead rate method, using direct machine hours as the allocation base for each department.

    Commercial $ per unit
    Residential $ per unit

    c. Recommend to management a product costing approach, based on your analyses in (a) and (b).

    1. The management should consider multiple production department factory overhead rate methods, because this method calculates the cost more accurately and considers the fact that commercial products use more costly overheads than residential products.

    2. The management should consider single plantwide factory overhead rate methods, because this method calculates the cost more accurately and considers the fact that the overheads are applied evenly based on the direct labor hours.

    3. The management could consider either multiple production department factory overhead rate method or the single plantwide rate, as both these methods have the same effect on the final costs.

    1
Answer #1

a.

Commercial $ 960 per Unit
Residential $ 540 per Unit

Calculation:

Plantwide overhead rate = 1,080,000 / 9,000 = 120

Per unit overhead allocated to Commercial = 120 x 8.0 = 960 per Unit

Per unit overhead allocated to Residential = 120 x 4.5 = 540 per Unit

b.

Commercial $1,100 per Unit
Residential $ 450 per Unit

Calculation:

Departmental overhead rate in Assembly = 280,000 / 4,000 = 70

Departmental overhead rate in Testing = 800,000 / 5,000 = 160

Costs Allocated to Product using Department overhead rate:

Commercial Residential
Assembly Department 140 (70*2) 210 (70*3)
Testing Department 960 (160*6) 240 (160*1.5)
Total Costs 1,100 450

c. The management should consider multiple production department factory overhead rate methods, because this method calculates the cost more accurately and considers the fact that commercial products use more costly overheads than residential products.

Explanation:

The reason for choosing multiple production department factory overhead rate method is that, both the departments have different budgeted overhead costs and different volumes of cost driver. So, the use of department rates will better show the correct cost of a product by taking into consideration the amount of cost driver used in each department. The allocation bases used for each department are likely to be more realistic in representing the relationship between overhead costs and the product, compared to using just one plantwide rate.